# Distributed Multi-Step Model Predictive Control for Consensus

### v1

@@ I-A Motivation and gap
Despite this progress, a recurring conceptual difficulty in DMPC consensus is the role of strictness. Moreau-type consensus theorems require, in one form or another, that when disagreement is present the next iterate enters the relative interior of a suitable neighbor hull, thereby preventing trajectories from remaining trapped on invariant faces of the global convex hull [7]. In many MPC-based protocols, convex-hull invariance can be enforced directly by terminal constraints, but strict interiority is often obtained indirectly via problem-specific geometric arguments (e.g., nonzero first step for integrators) or by strengthening terminal conditions in ways that may shrink feasibility sets and complicate robustness claims [3]. For general nonlinear dynamics and general convex constraints, it is therefore natural to ask: Which strictness property is truly necessary for consensus, and how can it be enforced by an MPC implementation without altering feasibility?
@@ I-B Contributions
Consensus as finite-step set stabilization. We recast consensus as stabilization of the agreement set and employ the disagreement diameter as a proper set-Lyapunov candidate. We restate a Moreau-type necessary/sufficient consensus theorem in our notation for time-varying directed graphs, highlighting that convex-hull inclusion and a strictness (relative-interior) property form the essential geometric ingredients.
A distributed multi-step MPC template enforcing hull inclusion and finite-step barycenter contraction. We introduce a local MPC problem that (a) constrains the terminal state to lie in the local neighbor convex hull, guaranteeing convex-hull invariance, and (b) enforces a finite-step contraction inequality toward a local barycenter. This yields a modular route to consensus via finite-step Lyapunov decay, provided that strictness can be ensured along closed-loop trajectories. We note that the feasibility of a finite-step Lyapunov decay is ensured by a constructive converse control finite-step Lyapunov function theorem, see [8] for more details.
Impossibility of deducing strictness from the baseline OCP and a minimal remedy. We show by an explicit counterexample that, even under feasibility and joint connectivity, the baseline OCP does not imply the required relative-interior strictness: feasible (and even cost-improving) solutions may lie on the boundary of the neighbor hull, allowing boundary-invariant behavior. To remedy this without shrinking the feasible set or changing the primary objective value, we introduce a two-stage lexicographic selection rule: among all (near-)optimal primary solutions, select one that maximizes an interiority measure (distance to the hull boundary). We prove that whenever an interior feasible terminal point exists, this tie-break enforces relative-interior updates and restores the strictness condition required by the geometric consensus theorem.
Explicit integrator specializations and computational evidence. For single- and double-integrator agents with input bounds (and no additional state constraints), we derive explicit horizon conditions guaranteeing feasibility and automatic existence of interior feasible terminal points, leading to consensus under the proposed lexicographic implementation. We provide distributed numerical simulations (including per-agent solve-time statistics and problem dimensions) demonstrating monotone decay of the diameter and illustrating when the lexicographic stage is activated in practice.
@@ III Problem Formulation
Assumption 1 (Basic regularity).
@@
$\|g_{i}(x_{i},u_{i})\|\leq\kappa_{1,i}(\|x_{i}\|)+\kappa_{2,i}(\|u_{i}\|),\qquad\forall x_{i},u_{i}.$
Assumption 1 ensures well-posedness of trajectories under bounded inputs and is standard in nonlinear MPC.
@@ III-B Finite-step control Lyapunov function and barycenter tracking
Outer sampling and information pattern.
@@
Let $M\in\mathbb{N}$ and $j\in\mathbb{Z}_{\geq 0}$ be given. Consider an agent (1) and a map $q_{i}:\mathcal{X}_{i}\times\bar{\mathcal{Z}}_{i}\to\mathcal{U}_{i}^{M}$. Partition $q_{i}=:(q_{i}^{1},\dots,q_{i}^{M})$, where $q_{i}^{k}\in\mathcal{U}_{i}$ for each $k=1,\dots,M$. The map $q_{i}$} is called an admissible finite-step feedback control law if for all $x_{i}(jM)\in\mathcal{X}_{i},\bar{z}_{i}(j)\in\bar{\mathcal{Z}}_{i}$ and all $k=1,\dots,M$, $g_{i}\big(x_{i}(jM+k-1),q_{i}^{k}(x_{i}(jM),\bar{z}_{i}(j))\big)\in\mathcal{X}_{i}$.
@@
Following the converse Lyapunov function [8, Theorem 10], under the decay condition $\beta(r,M)<r$ for $\beta\in\mathcal{KL}$ in (4), the existence of admissible finite-step feedback control law implies any scaling version of the norm function $\|\cdot\|$ itself is a so-called finite-step control Lyapunov function for the system converse control finite-step Lyapunov function theorem, see [8] for more details. This can be technically summarized as follows, which is a restatement of [8, Theorem 10] in our current context:
@@
The decay inequality (5) will be encoded later into our optimization control problem to enforce the state variables $x_{i}$ towards the local barycenter $\bar{z}_{i}$, see (12) below.
@@ IV Necessary and Sufficient Geometric Conditions for Consensus
In [7], Moreau presents a number of necessary and sufficient conditions for consensus for agents with nonlinear discrete-time dynamics. Towards this end, he provides a (strict) convexity condition, given by [7, Assumption 1] plus a number of graph connectivity conditions, see [7, Theorems 2, 3]. In this section, we reformulate these  [7, Assumption 1] to our wording context.
@@
Now we pack and restate [7, Theorem 1, Theorem 2] in the following theorem.
@@
Assumption 3 is the point-valued specialization of Moreau’s convexity and strictness assumptions (see [7, Assumption 1]), where our hull inclusion (10) corresponds to his convexity property and our RI strictness (11) corresponds to his strict inclusion property (expressed as RI membership when disagreement is nonzero). The sufficiency part of Theorem 1 is the directed-graph asymptotic-consensus result in [7, Theorem 2], restated in our notation. The necessity part stated here is a standard locality-based necessity: without (N1) the global hull may expand, and without (N2) the network can be split into groups without mutual influence, so agreement for all initial conditions is impossible.
@@ V Consensus via Distributed Multi-Step MPC
The barycenter-tracking term is a local stage objective encouraging movement toward local agreement. Motivated by the sufficiency condition of the approximate strictness (11), the optimization problem aims to minimize the distance to the barycenter. Motivated by the necessity of the convex hull inclusion (10), this condition is encoded as the terminal constraint in the local OCP.
@@ V-A Consensus under OCP (12) via enforced finite-step barycenter contraction
The following result is a direct consequence of the necessary condition (N1) in Theorem 1. This can be seen as a restatement of [7, Lemma 1].
@@ IX Numerical Simulations
In this section, we illustrate the performance of Algorithm 2 with lexicographic selection for agents with single- and double-integrator dynamics. The simulations are performed using a distributed Python implementation based on a REQ/REP communication architecture [6, 11], where the agents, the controllers, and the coordinator communicate. All simulations were performed on a Windows 11 (64-bit) machine equipped with a 13th Gen Intel Core i5-1335U processor (10 cores, 1.30 GHz base frequency) and 32 GB RAM. The implementation was carried out in Python 3.13. Each agent solves its local OCP in parallel, while a coordinator aggregates the open-loop control sequences and propagates the plant dynamics over the finite-step horizon.
@@ IX-A Simulation setup
We consider a network of $\ell=4$ agents in $\mathbb{R}^{2}$ with a time-invariant ring-form connected communication graph. The initial conditions are chosen such that the initial convex hull is non-singleton and $V(0)>0$.
For each outer step $j$, the agents solve their local OCP over a finite-step horizon $M$ as prescribed in Algorithm 2 with the lexicographic tolerance $\delta_{\mathrm{lex}}=10^{-5}$. We take the decay function $\alpha$ in the terminal inequality (12) linear as $\alpha_{i}(s)=0.8s$ and the control input bounds are chosen as $\|u_{i}\|\leq 1$ for all $i=1,\dots,4$.
@@
All figures reported below are generated directly from the distributed implementation of Algorithm 2.
@@ IX-B Single-integrator case
We first consider the single inetgrator dynamics $x_{i}(t+1)=x_{i}(t)+u_{i}(t)$ as in (19).
Figure 1 shows the evolution of the position states over the inner steps. The outer-step instants are marked explicitly. The trajectories converge asymptotically to a common consensus value.
@@
The corresponding control inputs are depicted in Figure 2. The control magnitudes remain bounded and decay as consensus is approached.
@@
The evolution of the disagreement diameter function $V(z(j))$ is shown in Figure 3. As predicted by Corollary 2, $V(x(j))$ decreases monotonically and converges to zero.
@@
The interiority measure $\phi_{i,j}$ for each agent is reported in Figure 4. For all outer steps where $\mathcal{C}_{i}(j)$ is non-singleton, we observe $\phi_{i,j}>0$, indicating that the terminal point lies in the relative interior of $\mathcal{C}_{i}(j)$.
@@
Consequently, the lexicographic stage of Algorithm 2 is not activated, as shown in Figure 5. The lexicographic flag remains zero throughout the simulation.
@@ IX-C Double-integrator case
The velocity trajectories are displayed in Figure 8. As predicted by Corollary 3, velocities converge to zero. The control inputs are shown in Figure 9.
@@
The diameter function $V_{r}(j)$ is depicted in Figure 10. Again, a monotonic decrease is observed, confirming asymptotic consensus.
@@
The interiority measure $\phi_{i,j}$ remains strictly positive whenever $\mathcal{C}_{i}(j)$ is non-singleton (Figure 11). Hence, the lexicographic stage remains inactive throughout the simulation (Figure 12).
@@
The decay of the primary cost $J_{i}^{\star}(j)$ further confirms the contraction mechanism induced by the terminal constraint, see Figure 13.
@@ IX-D Discussion
Thus, the lexicographic stage serves as a safeguard mechanism ensuring strict interiority when necessary, while remaining inactive when the primary problem already enforces the desired strictness property.
Overall, the distributed finite-step MPC with lexicographic selection achieves asymptotic consensus in both single- and double-integrator settings, fully consistent with the theoretical development of Section VIII.

### v2

@@ I-A Motivation and gap
Despite this progress, a recurring conceptual difficulty in DMPC consensus is the role of strictness. Moreau-type consensus theorems require, in one form or another, that when disagreement is present, the next iteration enters the relative interior of a suitable neighbor hull, thereby preventing trajectories from remaining trapped on invariant faces of the global convex hull [9]. In many MPC-based protocols, convex-hull invariance can be enforced directly by terminal constraints, but strict interiority is often obtained indirectly via problem-specific geometric arguments (e.g., nonzero first step for integrators) or by strengthening terminal conditions in ways that may shrink feasibility sets and complicate robustness claims [3]. For general nonlinear dynamics and general convex constraints, it is therefore natural to ask: Which strictness property is truly necessary for consensus, and how can it be enforced by an MPC implementation without altering feasibility?
@@ I-B Contributions
Consensus as finite-step set stabilization: We recast consensus as stabilization of the agreement set and employ the disagreement diameter as a proper set-Lyapunov candidate. We restate a Moreau-type necessary/sufficient consensus theorem in our notation for time-varying directed graphs, highlighting that convex-hull inclusion and a strictness (relative-interior) property form the essential geometric ingredients.
A distributed multi-step MPC template enforcing hull inclusion and finite-step barycenter contraction: We introduce a local MPC problem that (a) constrains the terminal state to lie in the local neighbor convex hull, guaranteeing convex-hull invariance, and (b) enforces a finite-step contraction inequality toward a local barycenter. This yields a modular route to consensus via finite-step Lyapunov decay [5], provided that strictness can be ensured along closed-loop trajectories. We note that the feasibility of a finite-step Lyapunov decay is ensured by a constructive converse control finite-step Lyapunov function theorem, see [10] for more details.
Impossibility of deducing strictness from the baseline OCP and a minimal remedy: We show by an explicit counterexample that, even under feasibility and joint connectivity, the baseline OCP does not imply the required relative-interior strictness: feasible (and even cost-improving) solutions may lie on the boundary of the neighbor hull, allowing boundary-invariant behavior. To remedy this without shrinking the feasible set or changing the primary objective value, we introduce a two-stage lexicographic selection rule: among all (near-)optimal primary solutions, select one that maximizes an interiority measure (distance to the hull boundary). We prove that whenever an interior feasible terminal point exists, this tie-break enforces relative-interior updates and restores the strictness condition required by the geometric consensus theorem.
Explicit integrator specializations and computational evidence: For single- and double-integrator agents with input bounds (and no additional state constraints), we derive explicit horizon conditions guaranteeing feasibility and automatic existence of interior feasible terminal points, leading to consensus under the proposed lexicographic implementation. We provide distributed numerical simulations (including per-agent solve-time statistics and problem dimensions) demonstrating monotone decay of the diameter and illustrating if the lexicographic stage is activated or not in practice.
@@ III Problem Formulation
Assumption 1 ($\mathcal{K}$-boundedness).
@@
$\|g_{i}(x_{i},u_{i})\|\leq\kappa_{1,i}(\|x_{i}\|)+\kappa_{2,i}(\|u_{i}\|),\qquad\forall(x_{i},u_{i})\in(\mathcal{X}_{i},\mathcal{U}_{i}).$
Assumption 1 ensures well-posedness of trajectories under bounded inputs and is standard in nonlinear MPC and nonlinear systems stability analysis [6, 5].
@@ III-B Finite-step control Lyapunov function and barycenter tracking
Outer sampling and information pattern
@@
Let $M\in\mathbb{N}$ and $j\in\mathbb{Z}_{\geq 0}$ be given. Consider an agent (1) satisfying Assumption 1 and a map $q_{i}:\mathcal{X}_{i}\times\bar{\mathcal{Z}}_{i}\to\mathcal{U}_{i}^{M}$. Partition $q_{i}=:(q_{i}^{1},\dots,q_{i}^{M})$, where $q_{i}^{k}\in\mathcal{U}_{i}$ for each $k=1,\dots,M$. The map $q_{i}$} is called an admissible finite-step feedback control law if for all $x_{i}(jM)\in\mathcal{X}_{i},\bar{z}_{i}(j)\in\bar{\mathcal{Z}}_{i}$ and all $k=1,\dots,M$, $g_{i}\big(x_{i}(jM+k-1),q_{i}^{k}(x_{i}(jM),\bar{z}_{i}(j))\big)\in\mathcal{X}_{i}$.
@@
Following the converse Lyapunov function [10, Theorem 10], under the decay condition $\beta(r,M)<r$ for $\beta\in\mathcal{KL}$ in (4), the existence of admissible finite-step feedback control law implies any scaling version of the norm function $\|\cdot\|$ itself is a so-called finite-step control Lyapunov function for the system converse control finite-step Lyapunov function theorem, see [10] for more details. This can be technically summarized as follows, which is a restatement of [10, Theorem 10] in our current context:
@@
The decay inequality (5) will be encoded later into our optimization control problem to enforce the state variables $x_{i}$ towards the local barycenter $\bar{z}_{i}$, see (12) below. Moreover, we note that the $\mathcal{K}$-boundedness condition as in Assumption 1 is necessary to conclude Proposition 1.
@@ IV Necessary and Sufficient Geometric Conditions for Consensus
In [9], Moreau presents a number of necessary and sufficient conditions for consensus for agents with nonlinear discrete-time dynamics. Towards this end, he provides a (strict) convexity condition, given by [9, Assumption 1] plus a number of graph connectivity conditions, see [9, Theorems 2, 3]. In this section, we reformulate these  [9, Assumption 1] to our wording context.
@@
Now we pack and restate [9, Theorem 1, Theorem 2] in the following theorem.
@@
Assumption 3 is the point-valued specialization of Moreau’s convexity and strictness assumptions (see [9, Assumption 1]), where our hull inclusion (10) corresponds to his convexity property and our RI strictness (11) corresponds to his strict inclusion property (expressed as RI membership when disagreement is nonzero). The sufficiency part of Theorem 1 is the directed-graph asymptotic-consensus result in [9, Theorem 2], restated in our notation. The necessity part stated here is a standard locality-based necessity: without (N1) the global hull may expand, and without (N2) the network can be split into groups without mutual influence, so agreement for all initial conditions is impossible.
@@ V Consensus via Distributed Multi-Step MPC
The barycenter-tracking term is a local stage objective encouraging movement toward local agreement. Motivated by the sufficiency condition of the approximate strictness (11), the optimization problem aims to minimize the distance to the barycenter. Motivated by the necessity of the convex hull inclusion (10), this condition is encoded as the terminal constraint in the local OCP. The feasibility of the inequality condition $\|x_{i}(M)-\bar{z}_{i}(j)\|\leq\alpha(\|x_{i}(0)-\bar{z}_{i}(j)\|)$ in (12) is ensured by Proposition 1 thanks to the assumption on the exitence of admissible finite-step feedback control laws $q_{i}$’s.
@@ V-A Consensus under OCP (12) via enforced finite-step barycenter contraction
The following result is a direct consequence of the necessary condition (N1) in Theorem 1. This can be seen as a restatement of [9, Lemma 1].
@@ IX Numerical Simulations
In this section, we illustrate the performance of Algorithm 2 with lexicographic selection for agents with single- and double-integrator dynamics. The simulations are performed using a distributed Python implementation based on a REQ/REP communication architecture [8, 13], where a coordinator facilitates communication among the agents, each local controller with the corresponding agent. All simulations were performed on a Windows 11 (64-bit) machine equipped with a 13th Gen Intel Core i5-1335U processor (10 cores, 1.30 GHz base frequency) and 32 GB RAM. The implementation was carried out in Python 3.13. Each agent solves its local OCP in parallel, while a coordinator aggregates the open-loop control sequences and propagates the agent dynamics over the finite-step horizon.
@@ IX-A Simulation setup
We consider a network of $\ell=4$ agents in $\mathbb{R}^{2}$ with a time-invariant ring-form connected communication graph. The initial conditions are chosen such that the initial convex hull is non-singleton and $V(0)>0$. For each outer step $j$, the agents solve their local OCP over a finite-step horizon $M$ as prescribed in Algorithm 2 with the lexicographic tolerance $\delta_{\mathrm{lex}}=10^{-5}$. We take the decay function $\alpha$ in the terminal inequality (12) linear as $\alpha_{i}(s)=0.8s$ and the control input bounds are chosen as $\|u_{i}\|\leq 1$ for all $i=1,\dots,4$.
@@ IX-B Single-integrator case
We first consider the single-inetgrator dynamics $x_{i}(t+1)=x_{i}(t)+u_{i}(t)$ as in (19). Figure 1 shows the evolution of the position states over the inner steps. The outer-step instants are marked explicitly. The trajectories converge asymptotically to a common consensus value. The corresponding control inputs are depicted in Figure 2. The control magnitudes remain bounded and decay as consensus is approached. The evolution of the disagreement diameter function $V(z(j))$ is shown in Figure 3. As predicted by Corollary 2, $V(x(j))$ decreases monotonically and converges to zero.
@@
The interiority measure $\phi_{i,j}$ for each agent is reported in Figure 4. For all outer steps where $\mathcal{C}_{i}(j)$ is non-singleton, we observe $\phi_{i,j}>0$, indicating that the terminal point lies in the relative interior of $\mathcal{C}_{i}(j)$. Consequently, the lexicographic stage of Algorithm 2 is not activated, as shown in Figure 5. The lexicographic flag remains zero throughout the simulation.
@@ IX-C Double-integrator case
The velocity trajectories are displayed in Figure 8. As predicted by Corollary 3, velocities converge to zero. The control inputs are shown in Figure 9. The diameter function $V_{r}(j)$ is depicted in Figure 10. Again, a monotonic decrease is observed, which confirms asymptotic consensus. The interiority measure $\phi_{i,j}$ remains strictly positive whenever $\mathcal{C}_{i}(j)$ is non-singleton (Figure 11). Hence, the lexicographic stage remains inactive throughout the simulation (Figure 12). The decay of the primary cost $J_{i}^{\star}(j)$ further confirms the contraction mechanism induced by the terminal constraint, see Figure 13.
@@ IX-D Discussion
Thus, the lexicographic stage serves as a safeguard mechanism ensuring strict interiority when necessary, while remaining inactive when the primary problem already enforces the desired strictness property. Overall, the distributed finite-step MPC with lexicographic selection achieves asymptotic consensus in both single- and double-integrator settings, fully consistent with the theoretical development of Section VIII.
@@ IX-E Computational complexity and implementation remarks
X Acknowledgment
The author sincerely thanks Lars Grüne and Fabian Wirth for insightful and motivating discussions that initiated and inspired this research work.